# Regression-Based Proximal Causal Inference

### v1

@@ 3 Application: Right Heart Catheterization Treatment Effect
To illustrate the proposed proximal two-stage regression approach in practice, we revisit the SUPPORT study. Recall that the objective is to estimate the treatment effect of right heart catheterization (RHC) performed within the first 24 hours of ICU stay on 30-day survival outcome. The SUPPORT dataset comprises detailed information on 5735 patients from a large prospective cohort study conducted at 5 medical centers.
As previously discussed, we proceed as in Tchetgen Tchetgen et al. (2023) and assign $Z$ = (pafi1, paco21) and $W$ = (ph1, hema1) [13]. Subsequently, the authors applied the proximal two-stage least-squares approach to estimate an effect of RHC on survival days equal to $\hat{\beta}_{a}(\text{Proximal})=-1.80$ with 95% C.I. $[-2.64,-0.96]$, such that on average a patient receiving RHC would live 1.8 days less than had he or she not received RHC. This result was significantly larger than that obtained by standard linear regression, adjusting for all covariates $\hat{\beta}(OLS)=-1.25$ days with 95% C.I. $[-1.80,-0.70]$.
@@
In our illustration, we proceed with the same proxy allocation as Tchetgen Tchetgen et al.[13]. To illustrate the method in the binary outcome case, we consider the dichotomized primary outcome indicating whether a patient survived at least 30 days. Furthermore, outcome confounding proxies (ph1, hema1) are likewise dichotomized according to their median value (0, lower than the median; 1, higher than the median) and then further combined into a polytomous variable $W$. For data analysis, following the proposed two-stage approach for the binary primary outcome, in the first stage, we fit a polytomous logistic regression for $W|A,Z,X$, under linear specification on the logit scale, i.e. $W\sim A+Z+X+Y$ from which we obtain fitted values $\sum_{k=1}^{3}\mathrm{logit}\widehat{\mathrm{Pr}}(W=k|A,Z,Y=1))$ which is then included as a covariate in the second stage logistic regression for $Y|A,Z,X$ as described in the previous Section.
$\mathrm{logit}\widehat{\mathrm{Pr}}(W=k|A,Z,Y,X))=\hat{\alpha}_{0k}^{*}+\hat{\alpha}_{ak}^{*}A+\hat{\alpha}_{zk}^{*}Z+\hat{\alpha}_{xk}^{*}X+\hat{\tilde{\alpha}}_{yk}Y,\ \text{where }k\in\{1,2,3\}$
$\mathrm{logit}\widehat{\mathrm{Pr}}(Y=1|A,Z,W,X))=\hat{\beta}_{0}^{*}+\hat{\beta}_{a}^{*}A+\hat{\beta}_{x}^{*}X+\hat{\beta}_{u}^{*}\sum_{k=1}^{3}\mathrm{logit}\widehat{\mathrm{Pr}}(W=k|A,Z,Y=1))+\sum_{k=1}^{3}\hat{\tilde{\beta}}_{wk}I(W=k),$
where $\text{where }\hat{\beta}_{a}^{*}=\hat{\beta}_{a}$
The approach yields an estimated treatment effect $\hat{\beta}_{a}(Proximal)=-0.40$ with 95% C.I. $[-0.56,-0.26]$, such that on average a 0.40 decrease in the log odds of 30-day survival is anticipated for patients receiving RHC. In comparison, a standard logistic regression model, which includes treatment and adjusts for all measured covariates gives a slight attenuation of the estimated causal effect $\hat{\beta}_{a}(MLE)=-0.36$ with 95% C.I. $[-0.51,-0.21]$.
$\mathrm{logit}[\widehat{\mathrm{Pr}}(Y=1|A,Z,W,X)]=\hat{\beta}_{0}+\hat{\beta}_{a}A+\hat{\beta}_{z}Z+\hat{\beta}_{x}X+\sum_{k=1}^{3}\hat{\beta}_{wk}I(W=k)$
@@ 4 Discussion
In this paper, we have reviewed proximal G-computation via proximal two-stage least-squares regression approach due to Tchetgen Tchetgen et al (2023), which posits linear structural models for both treatment confounding proxies $W$ and the primary outcome $Y$[13]. The approach delivers valid causal inferences in the presence of hidden confounders for which proxies are available. Extending the approach, we have established analogous two-stage regression approaches for structural causal models with either log, and logit links apply to either $W$ or $Y$. For all possible combinations of link functions for the primary outcome and outcome confounding proxies, we have formulated an appropriate corresponding two-stage regression procedure, thus making it possible for epidemiologists to use the proximal causal approach by fitting familiar generalized linear models.
@@ Appendix A Appendix
In Sections A.1-A.6, we provide details for proximal two-stage regression for all cases outlined in Table 1 including various combinations of different link functions. We also introduce more flexible model setups that allow for interaction terms in cases where the logit link is involved in the first stage, the second stage, or both. The formulation for polytomous regression is considered in A.3.3. We suppress measured covariates $X$ for simplicity, which is consistent with DAG (2) in Figure 1. Note that all formulations and corresponding estimation procedures accommodate DAG (1) in Figure 1 where there is no direct causal link from $W$ to $Y$ and from $Z$ to $A$.
The proofs center on the mathematical property of link functions used to define GLMs. Consistent with Section 2.2, we implicitly assume proxies are valid and thus $W$ and $Z$ are $U-$relevant such that the coefficients encoding the relationship between them are non-zero throughout the proofs. We use $\mathbf{M}(*)$ to denote the moment-generating function (MGF) with input ***. In addition, $\propto_{L}$ is used to denote the linear transformation, say $Y\propto_{L}X$ means $Y=aX+b$ with constants $a$ and $b$ where $a\neq 0$, and $\propto$ is used to denote the proportionality, say $Y\propto X$ means $Y=aX$ with a non-zero constant $a$.
@@
*** ${\beta}_{0}^{*}=a\beta_{0}+b$ ($a\neq 0$) The linearly transformed parameter
@@
A.1 Identity Links Combination
@@
W Data Type Y Data Type
Continuous (Identity Link)
Count (Log Link)
Binary (Logit Link)
Continuous (Identity Link) Linear $W\sim A+Z$
$S=E[W|A,Z]$
Linear $Y\sim A+S$
Linear $W\sim A+Z$
$S=E[W|A,Z]$
Poisson $Y\sim A+S$
Linear $W\sim A+Z+Y$
$S=E[W|A,Z,Y=1]$
Logistic $Y\sim A+S$
Count (Log Link)
Poisson $W\sim A+Z$
$S=\mathrm{log}(E[W|A,Z])$
Linear $Y\sim A+S$
Poisson $W\sim A+Z$
$S=\mathrm{log}(E[W|A,Z])$
Poisson $Y\sim A+S$
Poisson $W\sim A+Z+Y$
$S=\mathrm{log}(E[W|A,Z,Y=1])$
Logistic $Y\sim A+S$
Binary (Logit Link)
Logistic $W\sim A+Z$
$S=\mathrm{logit}(\mathrm{Pr}(W=1|A,Z))$
Linear $Y\sim A+S+W$
Logistic $W\sim A+Z$
$S=\mathrm{logit}(\mathrm{Pr}(W=1|A,Z))$
Poisson $Y\sim A+S+W$
Logistic $W\sim A+Z+Y$
$S=\mathrm{logit}(\mathrm{Pr}(W=1|A,Z,Y=1))$
Logistic $Y\sim A+S+W$
@@
Proximal control variable $S$ is a linear transformation of $E[U|A,Z]$, i.e.,
@@ A.2 Log Links Combination
W Data Type Y Data Type
Continuous (Identity Link)
Count (Log Link)
Binary (Logit Link)
Continuous (Identity Link)
Linear $W\sim A+Z$
$S=E[W|A,Z]$
Linear $Y\sim A+S$
Linear $W\sim A+Z$
$S=E[W|A,Z]$
Poisson $Y\sim A+S$
Linear $W\sim A+Z+Y$
$S=E[W|A,Z,Y=1]$
Logistic $Y\sim A+S$
Count (Log Link)
Poisson $W\sim A+Z$
$S=\mathrm{log}(E[W|A,Z])$
Linear $Y\sim A+S$
Poisson $W\sim A+Z$
$S=\mathrm{log}(E[W|A,Z])$
Poisson $Y\sim A+S$
Poisson $W\sim A+Z+Y$
$S=\mathrm{log}(E[W|A,Z,Y=1])$
Logistic $Y\sim A+S$
Binary (Logit Link)
Logistic $W\sim A+Z$
$S=\mathrm{logit}(\mathrm{Pr}(W=1|A,Z))$
Linear $Y\sim A+S+W$
Logistic $W\sim A+Z$
$S=\mathrm{logit}(\mathrm{Pr}(W=1|A,Z))$
Poisson $Y\sim A+S+W$
Logistic $W\sim A+Z+Y$
$S=\mathrm{logit}(\mathrm{Pr}(W=1|A,Z,Y=1))$
Logistic $Y\sim A+S+W$
@@
$*.\displaystyle|E[U|A,Z]|<\infty;\textrm{ $\mathbf{M}_{\epsilon}(*)$, the MGF of$\epsilon$, exists for any real value *.}$.
@@
=$\stackrel{{\scriptstyle(c)}}{{=}}exp\{\alpha_{0}+\alpha_{u}E[U|A,Z]+\mathrm{log}\mathbf{M}_{\epsilon}(\alpha_{u})$
@@
Proximal control variable $S$ is a linear transformation of $E[U|A,Z]$, i.e.,
@@
A.3 Logit Links Combination
In this section, we present details on the binary $Y$ and binary $W$ cases, which corresponds to the gray cell in the table below.
W Data Type Y Data Type
Continuous (Identity Link)
Count (Log Link)
Binary (Logit Link)
Continuous (Identity Link)
Linear $W\sim A+Z$
$S=E[W|A,Z]$
Linear $Y\sim A+S$
Linear $W\sim A+Z$
$S=E[W|A,Z]$
Poisson $Y\sim A+S$
Linear $W\sim A+Z+Y$
$S=E[W|A,Z,Y=1]$
Logistic $Y\sim A+S$
Count (Log Link)
Poisson $W\sim A+Z$
$S=\mathrm{log}(E[W|A,Z])$
Linear $Y\sim A+S$
Poisson $W\sim A+Z$
$S=\mathrm{log}(E[W|A,Z])$
Poisson $Y\sim A+S$
Poisson $W\sim A+Z+Y$
$S=\mathrm{log}(E[W|A,Z,Y=1])$
Logistic $Y\sim A+S$
Binary (Logit Link)
Logistic $W\sim A+Z$
$S=\mathrm{logit}(\mathrm{Pr}(W=1|A,Z))$
Linear $Y\sim A+S+W$
Logistic $W\sim A+Z$
$S=\mathrm{logit}(\mathrm{Pr}(W=1|A,Z))$
Poisson $Y\sim A+S+W$
Logistic $W\sim A+Z+Y$
$S=\mathrm{logit}(\mathrm{Pr}(W=1|A,Z,Y=1))$
Logistic $Y\sim A+S+W$
@@ A.3 Logit Links Combination
$\mathrm{logit}(\mathrm{Pr}(Y=1|A,Z))\neq E[\mathrm{logit}(\mathrm{Pr}(Y=1|A,Z))]=\beta_{0}^{*}+\beta_{u}E[Y|A,Z],$
$\mathrm{logit}(\mathrm{Pr}(W=1|A,Z))\neq E[\mathrm{logit}(\mathrm{Pr}(Y=1|A,Z))]=\alpha_{0}^{*}+\alpha_{u}E[U|A,Z].$
@@
$OR(A,B,C)$ $=\frac{f(A,B|C)\cdot f(A=0,B=0|C)}{f(A=0,B|C)\cdot f(A,B=0|C)}$
@@
A.3.1 Logit Links Combination with No Interaction Term
@@
$*.\displaystyle|E[U|A,Z,Y=0,W=0]|<\infty;\textrm{ $\mathbf{M}_{\epsilon}(*)$, the MGF of $\epsilon|Y=0,W=0$ exists for any real value *.}$.
@@
Equations 48 and 49 imply Result 3 of the main paper
@@
A.3.2 Logit Links Combination with Interaction Terms
@@
$*.\displaystyle|E[U|A,Z|Y=0,W=0]|<\infty;\textrm{ $\mathbf{M}_{\epsilon}(*)$, the MGF of$\epsilon|Y=0,W=0$ exists for any real value *.}$.
@@
Proximal control variable $S=\mathrm{logit}(\mathrm{Pr}(W=1|A,Z,Y=1))$ is a linear transformation of $E[U|A,Z,W=0,Y=0]=m(A,Z)$, i.e.,
@@
A.3.3 Extension to Polytomous Regression
We present the derivation for the logit links case in which $W$ has $K+1$ categories $(0,\ldots,K)$ and $Y$ has $(T+1)$ categories $(0,\ldots,T)$. We set 00 as the reference level for $W$ and $Y$ and proceed with the following parameterization:
@@ A.3.3 Extension to Polytomous Regression
$*.\displaystyle|E[U|A,Z|Y=0,W=0]|<\infty;\textrm{ $\mathbf{M}_{\epsilon}(*)$, the MGF of $\epsilon|Y=0,W=0$ exists for any real value *.}$.
@@
We now prove Result 4 under Assumption 5.
@@
$\propto f(W=k|A,Z,Y,U=0)f(U|A,Z,Y,W=0)OR(W=k,U|A,Z,Y)$
$\propto exp\{\alpha_{0k}+\sum_{t=1}^{T}\alpha_{ykt}I(Y=t)\}f(U|A,Z,Y,W=0)exp\{\alpha_{uk}U\}$
@@
$=exp\{\alpha_{0k}+\mathrm{log}(\mathbf{M}_{\epsilon}(\alpha_{uk}))+\alpha_{uk}m(A,Z)+\sum_{t=1}^{T}\left(\alpha_{ykt}+\mathrm{log}\left(\frac{\mathbf{M}_{\epsilon}(\alpha_{uk}+\beta_{ut})}{\mathbf{M}_{\epsilon}(\alpha_{uk})\mathbf{M}_{\epsilon}(\beta_{ut})}\right)I(Y=t)\right)\}$
$=exp\{\tilde{\alpha}_{0k}+\alpha_{uk}m(A,Z)+\sum_{t=1}^{T}\tilde{\alpha}_{ykt}I(Y=t)\}$ (68)
$\textrm{where }\tilde{\alpha}_{0k}=\alpha_{0k}+\mathrm{log}(\mathbf{M}_{\epsilon}(\alpha_{uk})),\ \tilde{\alpha}_{ykt}=\alpha_{ykt}+\mathrm{log}\left(\frac{\mathbf{M}_{\epsilon}(\alpha_{uk}+\beta_{ut})}{\mathbf{M}_{\epsilon}(\alpha_{uk})\mathbf{M}_{\epsilon}(\beta_{ut})}\right)$
Following analogous algebraic steps in the binary $(Y,W)$ case, the transition from the second line to the third line is evident. Note the complete proof entails the representation from the Bayes’ rules in the same way to work on $f(W=k,U|A,Z,Y=t)$ for each $t\in\{1,\ldots,T\}$, which is omitted here.
@@
$\propto exp\{\beta_{0t}+\mathrm{log}(\mathbf{M}_{\epsilon}(\beta_{ut}))+\beta_{at}A+\sum_{k=1}^{K}\left(\beta_{wtk}+\mathrm{log}\left(\frac{\mathbf{M}_{\epsilon}(\alpha_{uk}+\beta_{ut})}{\mathbf{M}_{\epsilon}(\alpha_{uk})\mathbf{M}_{\epsilon}(\beta_{ut})}\right)I(W=k)\right)\}$
$\propto exp\{\tilde{\beta}_{0t}+\beta_{at}A+\sum_{k=1}^{K}\tilde{\beta}_{wtk}I(W=k)\}$ (69)
$\textrm{where }\tilde{\beta}_{0t}=\beta_{0t}+\mathrm{log}(\mathbf{M}_{\epsilon}(\beta_{ut})),\ \tilde{\beta}_{wtk}=\beta_{wtk}+\mathrm{log}\left(\frac{\mathbf{M}_{\epsilon}(\alpha_{uk}+\beta_{ut})}{\mathbf{M}_{\epsilon}(\alpha_{uk})\mathbf{M}_{\epsilon}(\beta_{ut})}\right)$
@@
$f(W=k|A,Z,Y)=\frac{exp\{\tilde{\alpha}_{0k}+\alpha_{uk}m(A,Z)+\sum_{t=1}^{T}\tilde{\alpha}_{ykt}I(Y=t)\}}{1+\sum_{k=1}^{K}exp\{\tilde{\alpha}_{0k}+\alpha_{uk}m(A,Z)+\sum_{t=1}^{T}\tilde{\alpha}_{ykt}I(Y=t)\}}$ (70)
$f(Y=t|A,Z,W)=\frac{exp\{\tilde{\beta}_{0t}+\beta_{at}A+\beta_{ut}m(A,Z)+\sum_{k=1}^{K}\tilde{\beta}_{wtk}I(W=k)\}}{1+\sum_{t=1}^{T}exp\{\tilde{\beta}_{0t}+\beta_{at}A+\beta_{ut}m(A,Z)+\sum_{k=1}^{K}\tilde{\beta}_{wtk}I(W=k)\}}$ (71)
@@
$\mathrm{logit}(\mathrm{Pr}(W=k|A,Z,Y))=\tilde{\alpha}_{0k}+\alpha_{uk}m(A,Z)+\sum_{t=1}^{T}\tilde{\alpha}_{ykt}I(Y=t)$ (72)
$\mathrm{logit}(\mathrm{Pr}(Y=t|A,Z,W))=\tilde{\beta}_{0t}+\beta_{at}A+\beta_{ut}m(A,Z)+\sum_{k=1}^{K}\tilde{\beta}_{wtk}I(W=k)$ (73)
Equations (72) and (73) imply Result 4 of the main paper.

### v3

@@ 3 Application: Right Heart Catheterization Treatment Effect
To illustrate the proposed proximal two-stage regression approach in practice, we revisit the SUPPORT study. The objective is to estimate the treatment effect of right heart catheterization (RHC) performed within the first 24 hours of ICU stay on 30-day survival outcome. The SUPPORT dataset comprises information on 5735 patients from a large prospective cohort study conducted at 5 medical centers. The treatment $A$ indicates receiving RHC treatment, and the outcome $Y$ is the survival time up to 30 days. The study also collected rich information encoded in 73 covariates including basic demographics (e.g. age, sex, race, education, income and insurance status), estimated probability of survival, comorbidity, vital signs, physiological status and functional status. Ten covariates measuring a patient’s overall physiological status were measured from a blood test drawn within 24 hours in the ICU: serum sodium, serum potassium, serum creatinine, bilirubin, albumin, PaO2/(.01*FiO2) ratio, PaCO2, serum PH (arterial), white blood cell count, and hematocrit. These variables may be subject to substantial measurement error and as single snapshots of the underlying physiological state over time can be considered potential confounding proxies.
[…]
@@
[…]
$\mathrm{logit}(\widehat{\mathrm{Pr}}(W=k|A,Z,X,Y))=\hat{\alpha}_{0k}^{*}+\hat{\alpha}_{ak}^{*}A+\hat{\alpha}_{zk}^{*}Z+\hat{\alpha}_{xk}^{*}X+\hat{\tilde{\alpha}}_{yk}Y,\ \text{where }k\in\{1,2,3\},$
$\mathrm{logit}(\widehat{\mathrm{Pr}}(Y=1|A,Z,X,W))=\hat{\beta}_{0}^{*}+\hat{\beta}_{a}^{*}A+\hat{\beta}_{x}^{*}X+\hat{\beta}_{u}^{*}\sum_{k=1}^{3}\mathrm{logit}(\widehat{\mathrm{Pr}}(W=k|A,Z,Y=0))+\sum_{k=1}^{3}\hat{\tilde{\beta}}_{wk}I(W=k),$
where $\text{where }\hat{\beta}_{a}^{*}=\hat{\beta}_{a}.$
The approach yields an estimated treatment effect $\hat{\beta}_{a}(Proximal)=-0.40$ with 95% C.I. $[-0.56,-0.26]$, such that on average a 0.40 decrease in the log odds of 30-day survival is anticipated for patients receiving RHC.
In comparison, a standard logistic regression model, which includes treatment and adjusts for all measured covariates gives a slight attenuation of the estimated causal effect $\hat{\beta}_{a}(MLE)=-0.36$ with 95% C.I. $[-0.51,-0.21]$.
$\mathrm{logit}(\widehat{\mathrm{Pr}}(Y=1|A,Z,W,X))=\hat{\beta}_{0}+\hat{\beta}_{a}A+\hat{\beta}_{z}Z+\hat{\beta}_{x}X+\sum_{k=1}^{3}\hat{\beta}_{wk}I(W=k)$
@@ 4 Discussion
In this paper, we have reviewed proximal G-computation via proximal two-stage least-squares regression approach due to Tchetgen Tchetgen et al. (2023), which posits linear structural models for both outcome confounding proxies $W$ and the primary outcome $Y$[13]. The approach delivers valid causal inferences in the presence of hidden confounders for which proxies are available. Extending the approach, we have established analogous two-stage regression approaches for structural causal models with either log, and logit links apply to either $W$ or $Y$. For all possible combinations of canonical link functions for the primary outcome and outcome confounding proxies, we have formulated an appropriate two-stage regression estimator detailed in Section A.1 of the Appendix, thus making it possible for epidemiologists to use the proximal causal approach by fitting familiar generalized linear models for the most common settings encountered in practice.
@@ Appendix A Appendix
In Section A.1, we introduce a unified formulation of the proximal two-stage regression approach under different combinations of GLMs with different link functions for $Y$ and $W$, which allows for addressing a variety of scenarios, such as the case where $Y$ is continuous while $W$ is binary. In the review of negative control methods by Shi et al. (2020) [6], several applications using negative controls in epidemiological studies were discussed, including examples with $Y$ for offspring weight, BMI, and blood pressure, and $W$ for paternal smoking [24, 25]. In these scenarios, $Y$ may be encoded as a continuous variable, while $W$ is encoded as a binary variable.
In Sections A.2-A.7, we provide details for proximal two-stage regression for all cases outlined in the table of Section A.1, including each combination of three link functions. We also present more flexible model setups that allow for interaction terms in model specification in cases where the logit link is involved in the first stage, the second stage, or both. The formulation for polytomous regression is considered in A.4.3. We suppress measured covariates $X$ for simplicity in derivation, which is consistent with DAG (2) in Figure 1. Note that all formulations and corresponding estimation procedures accommodate DAG (1) in Figure 1 where there is no direct causal link from $W$ to $Y$ and from $Z$ to $A$.
The proofs center on the mathematical property of link functions used to define GLMs. Consistent with Section 2.2, we implicitly assume proxies are valid and thus $W$ and $Z$ are $U-$relevant such that the coefficients encoding the relationship between them are non-zero throughout the proofs. We use $\mathbf{M}(*)$ to denote the moment-generating function (MGF) with input *. In addition, $\propto_{L}$ is used to denote the linear transformation, say $Y\propto_{L}X$ means $Y=aX+b$ with constants $a$ and $b$ where $a\neq 0$, and $\propto$ is used to denote the proportionality, say $Y\propto X$ means $Y=aX$ with a non-zero constant $a$.
@@
* ${\beta}_{0}^{*}=a\beta_{0}+b$ ($a\neq 0$) The linearly transformed parameter
@@
A.1 Unification of Proximal inference for GLMs
Figure 2: Estimation procedures for each combination of three link functions. $S$ denotes proximal control variable for $U$.
Assuming linear regression functions for the first- and the second-stage regression equations, the table of Figure 2 summarizes estimation procedures for each combination of three link functions introduced in Sections A.2-A.7. Importantly, each instance delivers a procedure that produces a consistent estimator of the causal effect of interest, whether or not there is a direct causal link from either $Z$ to $A$ or from $W$ to $Y$ because either the corresponding model specification is not contingent upon the presence of these direct causal links or the same estimation procedure applies to the model specification accounting for the presence of these direct causal.
For simplicity, thus far the exposition has ignored measured confounders $X$ and interaction terms, we now describe the ease with which they can be incorporated into the proposed methodology, by simply including them as covariates in regressions at both stages. Specifically, suppose that $Y$ and $W$ are both continuous and therefore modeled using linear models compatible with DAG (3) which now also involves measured confounders $X$, thus the first-stage linear model may take the form:
while the second-stage may take the form:
A similar approach applies to all other cases under different combinations of link functions, including those involving logit links where $(Y,W)-(A,Z,X)$ interactions might be considered by implementing a slightly more different estimation procedure (see Sections A.4.2, A.6.2, A.6.4, A.7.2 and A.7.4).
For inference, the bootstrap may be convenient in practice; however, we also provide analytic expressions for the asymptotic variance of parameter estimates obtained via proximal two-stage regression which we have implemented in R package ”pci2s”. The proof of asymptotic normality and standard error computation are also provided in Section A.9.
A.2 Identity Links Combination
@@
Proximal control variable $S$ is a linear transformation of $E[U|A,Z]$, i.e.
@@ A.3 Log Links Combination
$\mathbf{M}_{\epsilon}(*)$, the MGF of $\epsilon$, exists for any real value *.
@@
=$\stackrel{{\scriptstyle(c)}}{{=}}exp\{\alpha_{0}+\alpha_{u}E[U|A,Z]+\mathrm{log}\mathbf{M}_{\epsilon}(\alpha_{u})\}$
@@
Proximal control variable $S$ is a linear transformation of $E[U|A,Z]$, i.e.
@@
A.4 Logit Links Combination
In this section, we present details on the binary $Y$ and binary $W$ case, which corresponds to the gray cell in the table below.
@@ A.4 Logit Links Combination
$\mathrm{logit}(\mathrm{Pr}(Y=1|A,Z))\neq E[\mathrm{logit}(\mathrm{Pr}(Y=1|A,Z,U))|A,Z]=\beta_{0}^{*}+\beta_{u}E[Y|A,Z],$
$\mathrm{logit}(\mathrm{Pr}(W=1|A,Z))\neq E[\mathrm{logit}(\mathrm{Pr}(W=1|A,Z,U))|A,Z]=\alpha_{0}^{*}+\alpha_{u}E[U|A,Z].$
@@
$OR(A,B|C)$ $=\frac{f(A,B|C)\cdot f(A=0,B=0|C)}{f(A=0,B|C)\cdot f(A,B=0|C)}$
@@
A.4.1 Logit Links Combination with No Interaction Term
@@
Equations 50 and 51 imply Result 3 of the main paper
@@
A.4.2 Logit Links Combination with Interaction Terms
@@
Proximal control variable $S=\mathrm{logit}(\mathrm{Pr}(W=1|A,Z,Y=1))$ is a linear transformation of $E[U|A,Z,W=0,Y=0]=m(A,Z)$, i.e.
@@
A.4.3 Extension to Polytomous Regression
In this section, we present details on the multinomial $Y$ and multinomial $W$ case, which corresponds to the gray cell in the table below.
In the main text, $U$ is assumed to be a scalar throughout. Here, we consider a more general scenario with multidimensional $U$. We present the derivation for the logit links case in which $W$ has $K+1$ categories $(0,\ldots,K)$ (which is of dimension $K$) and $Y$ has $T+1$ categories $(0,\ldots,T)$. We assume $Z$ has a dimension of $Q$ and $U$ spans $R^{V}$ dimensions with $V\leq min(K,Q)$ so that the rank condition for identification is satisfied [15]. We set 00 as the reference level for $W$ and $Y$ and proceed with the following parameterization:
@@ A.4.3 Extension to Polytomous Regression
$\mathrm{logit}(\mathrm{Pr}(Y=t|A,Z,W,U))=\beta_{0t}+\beta_{at}A+\beta_{ut}^{\prime}U+\sum_{k=1}^{K}\beta_{wtk}I(W=k))$ (67)
$\mathrm{logit}(\mathrm{Pr}(W=k|A,Z,Y,U))=\alpha_{0k}+\alpha_{uk}^{\prime}U+\sum_{t=1}^{T}\alpha_{ykt}I(Y=t)$ (68)
@@
We note that $\beta_{ut},\alpha_{uk},m(A,Z)\in R^{V}$ of the same dimension of the vector $U=(U_{1},\ldots,U_{V})^{\prime}$
We now prove the general version of Result 4 under Assumption 5.
@@
$\propto f(W=k|A,Z,Y,U=\underline{0})f(U|A,Z,Y,W=0)OR(W=k,U|A,Z,Y)$
$\propto exp\{\alpha_{0k}+\sum_{t=1}^{T}\alpha_{ykt}I(Y=t)\}f(U|A,Z,Y,W=0)exp\{\alpha_{uk}^{\prime}U\}$
@@
$=exp\{\alpha_{0k}+\mathrm{log}(\mathbf{M}_{\epsilon}(\alpha_{uk}^{\prime}))+\alpha_{uk}^{\prime}m(A,Z)+\sum_{t=1}^{T}\left(\alpha_{ykt}+\mathrm{log}\left(\frac{\mathbf{M}_{\epsilon}(\alpha_{uk}^{\prime}+\beta_{ut}^{\prime})}{\mathbf{M}_{\epsilon}(\alpha_{uk}^{\prime})\mathbf{M}_{\epsilon}(\beta_{ut}^{\prime})}\right)\right)I(Y=t)\}$
$=exp\{\tilde{\alpha}_{0k}+\alpha_{uk}^{\prime}m(A,Z)+\sum_{t=1}^{T}\tilde{\alpha}_{ykt}I(Y=t)\}$ (70)
$\textrm{where }\tilde{\alpha}_{0k}=\alpha_{0k}+\mathrm{log}(\mathbf{M}_{\epsilon}(\alpha_{uk}^{\prime})),\ \tilde{\alpha}_{ykt}=\alpha_{ykt}+\mathrm{log}\left(\frac{\mathbf{M}_{\epsilon}(\alpha_{uk}^{\prime}+\beta_{ut}^{\prime})}{\mathbf{M}_{\epsilon}(\alpha_{uk}^{\prime})\mathbf{M}_{\epsilon}(\beta_{ut}^{\prime})}\right)$
Following analogous algebraic steps as in the binary $(Y,W)$ case, the transition from the second line to the third line becomes evident. Note the complete proof entails the representation from Bayes’ rule applied to $f(W=k,U|A,Z,Y=t)$ for each $t\in\{1,\ldots,T\}$, which is omitted here.
@@
$\propto exp\{\beta_{0t}+\mathrm{log}(\mathbf{M}_{\epsilon}(\beta_{ut}^{\prime}))+\beta_{at}A+\beta_{ut}^{\prime}m(A,Z)+\sum_{k=1}^{K}\left(\beta_{wtk}+\mathrm{log}\left(\frac{\mathbf{M}_{\epsilon}(\alpha_{uk}^{\prime}+\beta_{ut}^{\prime})}{\mathbf{M}_{\epsilon}(\alpha_{uk}^{\prime})\mathbf{M}_{\epsilon}(\beta_{ut}^{\prime})}\right)\right)I(W=k)\}$
$\propto exp\{\tilde{\beta}_{0t}+\beta_{at}A+\beta_{ut}^{\prime}m(A,Z)+\sum_{k=1}^{K}\tilde{\beta}_{wtk}I(W=k)\}$ (71)
$\textrm{where }\tilde{\beta}_{0t}=\beta_{0t}+\mathrm{log}(\mathbf{M}_{\epsilon}(\beta_{ut}^{\prime})),\ \tilde{\beta}_{wtk}=\beta_{wtk}+\mathrm{log}\left(\frac{\mathbf{M}_{\epsilon}(\alpha_{uk}^{\prime}+\beta_{ut}^{\prime})}{\mathbf{M}_{\epsilon}(\alpha_{uk}^{\prime})\mathbf{M}_{\epsilon}(\beta_{ut}^{\prime})}\right)$
@@
$f(W=k|A,Z,Y)=\frac{exp\{\tilde{\alpha}_{0k}+\alpha_{uk}^{\prime}m(A,Z)+\sum_{t=1}^{T}\tilde{\alpha}_{ykt}I(Y=t)\}}{1+\sum_{k=1}^{K}exp\{\tilde{\alpha}_{0k}+\alpha_{uk}^{\prime}m(A,Z)+\sum_{t=1}^{T}\tilde{\alpha}_{ykt}I(Y=t)\}}$ (72)
$f(Y=t|A,Z,W)=\frac{exp\{\tilde{\beta}_{0t}+\beta_{at}A+\beta_{ut}^{\prime}m(A,Z)+\sum_{k=1}^{K}\tilde{\beta}_{wtk}I(W=k)\}}{1+\sum_{t=1}^{T}exp\{\tilde{\beta}_{0t}+\beta_{at}A+\beta_{ut}^{\prime}m(A,Z)+\sum_{k=1}^{K}\tilde{\beta}_{wtk}I(W=k)\}}$ (73)
@@
$\mathrm{logit}(\mathrm{Pr}(W=k|A,Z,Y))=\tilde{\alpha}_{0k}+\alpha_{uk}^{\prime}m(A,Z)+\sum_{t=1}^{T}\tilde{\alpha}_{ykt}I(Y=t)$ (74)
$\mathrm{logit}(\mathrm{Pr}(Y=t|A,Z,W))=\tilde{\beta}_{0t}+\beta_{at}A+\beta_{ut}^{\prime}m(A,Z)+\sum_{k=1}^{K}\tilde{\beta}_{wtk}I(W=k)$ (75)
Equations (74) and (75) imply a more general version of Result 4 of the main paper.